OpenStudy (anonymous):

I WILL FAN AND MEDAL!!! PLZ HELP??? Marks: 2 Which table of values reflects a direct variation? Choose one answer. A. A table shows that when x is zero, y is zero; when x is 1, y is 5; and when x is 2, y is 6. B. A table shows that when x is 1, y is 5; when x is 2, y is 5; and when x is 3, y is 5. C. A table shows that when x is 2, y is 4; when x is 4, y is sixteen; and when x is 6, y is thirty-six. D. A table shows that when x is zero, y is zero; when x is 3, y is 6; and when x is 5, y is ten.

OpenStudy (anonymous):

@Kash_TheSmartGuy @gina123

OpenStudy (kash_thesmartguy):

Mostly, I would go with C.

OpenStudy (kash_thesmartguy):

Or D.

OpenStudy (kash_thesmartguy):

Mostly, C.

OpenStudy (johnweldon1993):

 So basically a direct variation is when \(\large y = kx\) where 'k' is some constant of variation we need to find Once we find 'k' we can use it to see if the direct variation holds true for the other numbers in question as well

OpenStudy (anonymous):

D? Because y = kx y = 2x y = 2(0) = 0 y = 2(3) = 6 y = 2(5) = 10

OpenStudy (kash_thesmartguy):

Yes, it's D. Confirm it.

OpenStudy (kash_thesmartguy):

I checked it out and got D.

OpenStudy (anonymous):

I has the same constant. All the others do not have a constant, but rather are quadratic or exponential in shape.

OpenStudy (anonymous):

ok thx